the length of a rectangular garden is 6 inches more than its width this area of the garden is 91 square inches find the dimensions of the garden
Answer by addingup(3677)

Length x Width = 91
Length = W+6 substitute in the formula above:
(W+6) x W = 91 distributive properties of multiplication: distribute the W:
W x W + 6 x W = W^2+6W = 91
Now we have two options to solve this equation, we can either complete the squares or use the quadratic formula. I'm going to complete the squares, if you need the quadratic formula send me a note.
W^2+6W = 91
Add 9 to both sides:
W^2+6W+9 = 100
Write the left hand side as a square:
(W+3)^2 = 100
Now we just take the square root of both sides and we're almost done:
W+3 = 10 or W+3 = -10
W = 7 or W = -13
Since we are not looking for a negative number, let's try the 7 in the original equation:
W = 7
L = W+6 = 7+6 = 13
L x W = 91
13 x 7 = 91 Correct.
:
Happy learning,
John